OpenStudy (anonymous):

What are the solutions of 4x2 = −20x? x = 0, x = 5 x = 0, x = −5 x = 4, x = 5 x = 4, x = −5

OpenStudy (unklerhaukus):

if you substitute x= a solution the equality should be true,

OpenStudy (anonymous):

the last option i think

OpenStudy (unklerhaukus):

\[4x^2 = −20x\] lets test x=4 does \[4(4)^2 = −20(4)\] \[64=-80\] this is not true so x=4 is not a solution